# Improving the Accuracy of DC Optimal Power Flow Formulations via Parameter Optimization

### v1

@@ II-B The DC-OPF Problem
The DC-OPF problem is a linear approximation of the AC-OPF problem that assumes uniform voltage magnitudes (typically normalized to one per-unit), small voltage angle differences, negligible losses, and ignores reactive power [19]. These assumptions are most relevant to transmission systems. The DC-OPF model underlies various applications within electricity markets and is widely used in unit commitment, transmission network expansion planning, and optimal transmission switching problems [9, 26, 27, 28, 29, 22].
@@
$\min_{p^{\text{g}}_{i},\theta_{i}}\quad$ $\sum_{i\in\mathcal{N}}c_{2i}(p^{\text{g}}_{i})^{2}+c_{1i}p^{\text{g}}_{i}+c_{0i}$ (2a)
@@
$p^{\text{g}}_{i}-\text{p}^{\text{d}}_{i}-\gamma_{i}=\sum_{(i,j)\in\mathcal{E}}P_{ij}+\sum_{(k,i)\in\mathcal{E}}P_{ki}$ (2d)
$P_{jk}=b_{jk}(\theta_{k}-\theta_{j})+\rho_{jk}$ (2e)
@@
[…]
Model 2 presents the DC-OPF problem. Analogous to (1a), the objective (2a) minimizes the total generation cost by summing quadratic functions of the generators’ active power outputs. Similarly, (2b) bounds the generators’ active power outputs, reminiscent of the dispatch bounds (1c). Analogous to (1d), constraint (2c) enforces thermal limits on transmission lines. Analogous to (1e), constraint (2d) balances active power at each bus by equating the net power injection with the net branch power flow. Constraint (2e) linearizes the line flow expressions by relating branch power flows to voltage angle differences according to the DC power flow approximation, thus offering a simplified counterpart to (1f) and (1g). Finally, (2f) sets the reference bus angle to zero. Note that (2d) and (2e) include parameters $\bm{\gamma}$ and $\bm{\rho}$. The parameter $\gamma_{i}$ accounts for losses from shunts, HVDC infeeds, and injections modeling phase shifts and branch losses for lines connected to bus $i$. The parameter $\rho_{ij}$ account for losses in line $(i,j)\in\mathcal{E}$.
Model 3 equivalently rewrites Model 2 in a more notationally friendly matrix form. In Model 3, we define $\mathbf{A}$ as the ||×$|\mathcal{E}|\times|\mathcal{N}|$ incidence matrix describing the connections between the system’s buses and branches and $\text{diag}(\mathbf{b})$ as a diagonal matrix having the entries of $\mathbf{b}$ on its diagonal.
@@
$\min_{\mathbf{p}^{\text{g}},\bm{\theta}}\quad$ ${\mathbf{p}^{\text{g}}}^{\top}\text{diag}(\mathbf{c}_{2})\mathbf{p}^{\text{g}}+\mathbf{c}^{\top}_{1}\mathbf{p}^{\text{g}}+\sum_{i\in\mathcal{N}}c_{0i}$ (3a)
@@
[…]
The DC-OPF problem’s accuracy relative to the AC-OPF problem is dictated by the parameters $\mathbf{b}$, $\bm{\gamma}$, and $\bm{\rho}$. Traditionally, there are two DC-OPF versions, cold-start and hot-start, that rely on different amounts of prior information to select these parameters.
@@ II-B1 Cold-start DC-OPF
$b^{cold}_{ij}=\Im\left(\frac{-1}{r_{ij}+\mathbf{j}x_{ij}}\right).$ (4)
@@ II-B2 Hot-start DC-OPF
Hot-start DC-OPF problems leverage information from a nominal AC power flow solution to improve accuracy. There are many hot-start DC-OPF variants [19]; for example, the “localized loss modeling” variant of the hot-start DC power flow model in [19] selects the coefficient and bias parameters as:
$b^{hot}_{ij}$ $=b_{ij}v^{\bullet}_{i}v^{\bullet}_{j}\sin(\theta^{\bullet}_{i}-\theta^{\bullet}_{j})/({\theta^{\bullet}_{i}-\theta^{\bullet}_{j}}),$ (5a)
$\gamma^{hot}_{i}$ $=\sum_{(i,j)\in\mathcal{E}}\Re(Y_{ij})v^{\bullet}_{i}(v^{\bullet}_{i}-v^{\bullet}_{j}\cos(\theta^{\bullet}_{i}-\theta^{\bullet}_{j})),$ (5b)
$\rho^{hot}_{ij}$ $=\Re(Y_{ij})v^{\bullet}_{i}(v^{\bullet}_{i}-v^{\bullet}_{j}\cos(\theta^{\bullet}_{i}-\theta^{\bullet}_{j})),$ (5c)
where the notation $(\,\cdot\,)^{\bullet}$ signifies quantities associated with a nominal AC power flow solution. The bias $\bm{\gamma}^{hot}$ aims to capture the effects of branch losses on phase angles and $\bm{\rho}^{hot}$ incorporates the branch losses in the line flow expressions.
II-B3 Optimized Parameters for DC Power Flow (DCPF) [23]
@@
Our recent work in [23] proposes an alternative for selecting the $\mathbf{b}$, $\bm{\gamma}$, and $\bm{\rho}$ parameters by solving an optimization problem. Providing a conceptual foundation for the approach we will propose in this paper, the algorithm in [23] also optimizes these parameters over a range of operation. However, there is a key distinction between the algorithm in [23] and the parameter optimization algorithm in this paper. Specifically, the approach in [23] does not explicitly consider the impacts of the optimized parameters on the solution to a DC-OPF problem, but rather seeks to balance the DC power flow’s accuracy with respect to all operating points within a specified range. Conversely, our proposed algorithm in this paper leverages machine learning techniques for training models with differentiable optimization layers to tailor the DC power flow parameter values so that the DC optimal power flow solutions align with the AC-OPF solutions. As shown numerically in Section IV, this enables further accuracy improvements relative to [23] in settings where the optimized DC power flow parameters are used in optimization problems like DC-OPF.
@@
We next introduce our machine learning-inspired algorithm for optimizing the coefficient ($\mathbf{b}$) and bias ($\bm{\gamma}$ and $\bm{\rho}$) parameters. Our algorithm seeks to match the generators’ active power outputs from the DC-OPF solution to corresponding values from the AC-OPF problem.
@@ III Parameter Optimization Algorithm
Our parameter optimization algorithm, depicted in Fig. 1, has offline and online phases. The offline phase computes DC-OPF parameters ($\mathbf{b}$, $\bm{\gamma}$, and $\bm{\rho}$) to achieve a high-fidelity approximation of the AC-OPF solution over a range of operational conditions. The online phase employs these parameters to efficiently solve DC-OPF problems in various applications. While our numerical results focus on DC-OPF problems, the optimized DC-OPF parameters are applicable to many computationally intensive optimization problems that incorporate DC power flow models, such as unit commitment [9, 26], expansion planning [3], and grid reliability enhancement [30].
@@
Our algorithm minimizes a loss function that quantifies the discrepancy between DC-OPF and AC-OPF generator active power setpoints over various load scenarios. To perform this minimization, we compute the loss function’s gradients with respect to the parameters $\mathbf{b}$, $\bm{\gamma}$, and $\bm{\rho}$ for use in a nonlinear optimization method. Based on our numerical tests comparing various nonlinear optimization methods, the TNC optimization method best used these gradients to achieve scalable and effective parameter optimization. TNC’s capability to approximate the Hessian matrix from gradient information ensures efficient handling of large-scale power systems.
@@ III-A Loss Function
We first introduce a loss function $\mathcal{L}$ that is the sum of squared two-norm discrepancies between the generators’ active power setpoints obtained from the AC-OPF problems ($\mathbf{p}_{m}^{\text{g,AC}}$) and the DC-OPF problems ($\mathbf{p}_{m}^{\text{g}}$) across a specified set of load scenarios $\mathcal{M}$:
$\mathcal{L}(\mathbf{b},\bm{\gamma},\bm{\rho})$ $=\frac{1}{|\mathcal{G}|\times|\mathcal{M}|}\sum_{m\in\mathcal{M}}||\mathbf{p}_{m}^{\text{g}}(\mathbf{b},\bm{\gamma},\bm{\rho})-\mathbf{p}_{m}^{\text{g,AC}}||^{2}_{2},$ (6)
where the constant $\frac{1}{|\mathcal{G}|\times|\mathcal{M}|}$ normalizes this function based on the number of generators and load scenarios. As shown in (6), $\mathbf{p}_{m}^{\text{g}}$ (and thus $\mathcal{L}(\mathbf{b},\bm{\gamma},\bm{\rho})$) is a function of the coefficient parameters $\mathbf{b}$ and the bias parameters $\bm{\gamma}$ and $\bm{\rho}$.
@@
$\min_{\mathbf{b},\bm{\gamma},\bm{\rho}}\quad\mathcal{L}(\mathbf{b},\bm{\gamma},\bm{\rho}).$ (7)
Using the sensitivity analysis which we describe in the next subsection, we adjust the coefficient and bias parameters ($\mathbf{b}$, $\bm{\gamma}$, and $\bm{\rho}$) to optimize the DC-OPF problem’s parameters. As detailed in Model 4, this task takes the form of a bilevel optimization problem, with the upper-level fine-tuning parameters based on outcomes from lower-level DC-OPF problems under a variety of operational scenarios.
@@
$\min_{\mathbf{b},\bm{\gamma},\bm{\rho}}\quad$ $\mathcal{L}=\frac{1}{|\mathcal{G}|\times|\mathcal{M|}}\sum_{m\in\mathcal{M}}||\mathbf{p}_{m}^{\text{g}}(\mathbf{b},\bm{\gamma},\bm{\rho})-\mathbf{p}^{\text{g},\text{AC}}_{m}||^{2}_{2}$ (8a)
@@
$\min_{\mathbf{p}^{\text{g}}_{m},\bm{\theta}_{m}}\quad{\mathbf{p}^{\text{g}}_{m}}^{\top}\text{diag}(\mathbf{c}_{2})\mathbf{p}^{\text{g}}_{m}+\mathbf{c}^{\top}_{1}\mathbf{p}^{\text{g}}_{m}+\sum_{i\in\mathcal{G}}c_{0i}$ (8b)
@@
$\quad\quad\mathbf{p}^{\text{g}}_{m}-\textbf{p}^{\text{d}}_{m}-\bm{\gamma}=\mathbf{A}^{\top}\Big{(}\text{diag}(\mathbf{b})\mathbf{A}\bm{\theta}_{m}+\bm{\rho}\Big{)}$ (8c)
@@
          $\alpha_{2}$: Curvature condition constant, between $\alpha_{1}$ and $1$
@@
[…]
@@ III-B Sensitivity Analysis of Coefficient and Bias Parameters
$\mathbf{g}^{b}=\frac{\partial\mathcal{L}}{\partial\mathbf{b}}=\frac{2}{|\mathcal{G}|\times|\mathcal{M}|}\sum_{m\in\mathcal{M}}\left.\frac{\partial\mathbf{p}^{\text{g}}}{\partial\mathbf{b}}\right|_{\mathbf{p}^{\text{g}}_{m}}\Big{(}\mathbf{p}_{m}^{\text{g}}-\mathbf{p}_{m}^{\text{g},AC}\Big{)},$ (9a)
$\mathbf{g}^{\gamma}=\frac{\partial\mathcal{L}}{\partial\bm{\gamma}}=\frac{2}{|\mathcal{G}|\times|\mathcal{M}|}\sum_{m\in\mathcal{M}}\left.\frac{\partial\mathbf{p}^{\text{g}}}{\partial\bm{\gamma}}\right|_{\mathbf{p}^{\text{g}}_{m}}\Big{(}\mathbf{p}_{m}^{\text{g}}-\mathbf{p}_{m}^{\text{g},AC}\Big{)},$ (9b)
$\mathbf{g}^{\rho}=\frac{\partial\mathcal{L}}{\partial\bm{\rho}}=\frac{2}{|\mathcal{G}|\times|\mathcal{M}|}\sum_{m\in\mathcal{M}}\left.\frac{\partial\mathbf{p}^{\text{g}}}{\partial\bm{\rho}}\right|_{\mathbf{p}^{\text{g}}_{m}}\Big{(}\mathbf{p}_{m}^{\text{g}}-\mathbf{p}_{m}^{\text{g},AC}\Big{)}.$ (9c)
The derivatives $\frac{\partial\mathbf{p}^{\text{g}}}{\partial\mathbf{b}}$, $\frac{\partial\mathbf{p}^{\text{g}}}{\partial\bm{\gamma}}$, and $\frac{\partial\mathbf{p}^{\text{g}}}{\partial\bm{\rho}}$ are computed from the KKT conditions of the lower-level DC-OPF problems (8b)–(8f) in Model 4. Differentiating through convex optimization problems like DC-OPF involves implicit differentiation of their optimality conditions [31, 32, 24]. To perform these implicit derivative computations, we leverage techniques developed for embedding convex optimization layers within deep learning architectures; specifically, we utilize cvxpylayers[25], which allows for automatic differentiation through convex optimization problems. This integration facilitates efficient and scalable gradient computations necessary for our optimization framework. In detail, we define the DC-OPF problem using CVXPY’s modeling syntax, specifying the objective function and constraints as per Model 3. This CVXPY model is then converted into a differentiable layer using cvxpylayers, enabling it to be embedded within our optimization pipeline. During the forward pass, cvxpylayer solves the DC-OPF problem to obtain $\mathbf{p}^{\text{g}}$. In the backward pass, cvxpylayer automatically computes the derivatives $\frac{\partial\mathbf{p}^{\text{g}}}{\partial\mathbf{b}}$, $\frac{\partial\mathbf{p}^{\text{g}}}{\partial\bm{\gamma}}$, and $\frac{\partial\mathbf{p}^{\text{g}}}{\partial\bm{\rho}}$, leveraging automatic differentiation provided by cvxpylayers; we refer the reader to [25, 24] for further details and efficient computational implementations.

### v3

@@ II-B The DC-OPF Problem
The DC-OPF problem is a linear approximation of the AC-OPF problem that assumes uniform voltage magnitudes (typically normalized to one per-unit), small voltage angle differences, negligible losses, and ignores reactive power [stott2009dc]. These assumptions are most relevant to transmission systems. The DC-OPF model underlies various applications within electricity markets and is widely used in unit commitment, transmission network expansion planning, and optimal transmission switching problems [vanAckooij2018, knueven2020mixed, fisher2008optimal, hedman2008optimal, guo2022tightening, Khodaei2010].
@@
$\operatorname*{minimize}_{p^{\text{g}}_{i},\theta_{i}}\quad$ $\sum_{i\in\mathcal{N}}c_{2i}(p^{\text{g}}_{i})^{2}+c_{1i}p^{\text{g}}_{i}+c_{0i}$ (2a)
@@
$p^{\text{g}}_{i}-\text{p}^{\text{d}}_{i}-\gamma_{i}=\sum_{(i,j)\in\mathcal{E}}P_{ij}-\sum_{(k,i)\in\mathcal{E}}P_{ki}$ (2d)
$P_{jk}=b_{jk}(\theta_{j}-\theta_{k})+\rho_{jk}$ (2e)
@@
$\textbf{variables}:$ $~p^{\text{g}}_{i}~(\forall i\in\mathcal{N}),\theta_{i}~(\forall i\in\mathcal{N})$
Model 2 presents the DC-OPF problem. Analogous to (1), (2a) minimizes the total generation cost, (2b) bounds the generators’ active power outputs, (2c) enforces line flow limits, and (2d) balances active power at each bus. Constraint (2e) linearizes the line flow expressions by relating branch power flows to voltage angle differences according to the DC power flow approximation, thus offering a simplified counterpart to (1f) and (1g). Finally, (2f) sets the reference angle.
Note that (2d) and (2e) include parameters $\bm{\gamma}$ and $\bm{\rho}$ to improve accuracy. Following extended DC power flow models [stott2009dc], these parameters have distinct roles. The flow bias $\rho_{ij}$ in (2e) can account for factors like phase-shifter effects and active power losses on the line $(i,j)$. The injection bias $\gamma_{i}$ in (2d) can account for losses from shunts, HVDC infeeds, or aggregate the effects of losses from all connected lines.
Model 3 equivalently rewrites Model 2 in a matrix form, where $\mathbf{A}$ is the $|\mathcal{E}|\times|\mathcal{N}|$ branch-bus incidence matrix and $\text{diag}(\mathbf{b})$ is the diagonal matrix with $\mathbf{b}$ on the diagonal.
@@
$\operatorname*{minimize}_{\mathbf{p}^{\text{g}},\bm{\theta}}\quad$ ${\mathbf{p}^{\text{g}}}^{\top}\text{diag}(\mathbf{c}_{2})\mathbf{p}^{\text{g}}+\mathbf{c}^{\top}_{1}\mathbf{p}^{\text{g}}+\mathbf{1}^{\top}\mathbf{c}_{0}$ (3a)
@@
$\textbf{variables}:$ $~\mathbf{p}^{\text{g}},\bm{\theta}$
The DC-OPF problem’s accuracy is dictated by the parameters $\mathbf{b}$, $\bm{\gamma}$, and $\bm{\rho}$. Traditionally, there are two DC-OPF versions, cold-start and hot-start, that rely on different amounts of prior information to select these parameters.
@@ II-B1 Cold-start DC-OPF
$b^{cold}_{ij}=\Im\left(\frac{-1}{r_{ij}+\mathrm{j}x_{ij}}\right).$ (4)
@@ II-B2 Hot-start DC-OPF
Hot-start DC-OPF problems leverage information from a nominal AC power flow solution to improve accuracy. There are many hot-start DC-OPF variants [stott2009dc]; e.g., the “localized loss modeling” variant in [stott2009dc] selects the coefficient and bias parameters as:
$b^{\text{hot}}_{ij}$ $=b_{ij}\,|V^{\bullet}_{i}|\,|V^{\bullet}_{j}|\frac{\sin(\theta^{\bullet}_{i}-\theta^{\bullet}_{j})}{\theta^{\bullet}_{i}-\theta^{\bullet}_{j}},$ (5a)
$\gamma^{hot}_{i}$ $=\sum_{(i,j)\in\mathcal{E}}\Re(Y_{ij})|V^{\bullet}_{i}|(|V^{\bullet}_{i}|-|V^{\bullet}_{j}|\cos(\theta^{\bullet}_{i}-\theta^{\bullet}_{j})),$ (5b)
$\rho^{hot}_{ij}$ $=\Re(Y_{ij})|V^{\bullet}_{i}|(|V^{\bullet}_{i}|-|V^{\bullet}_{j}|\cos(\theta^{\bullet}_{i}-\theta^{\bullet}_{j})),$ (5c)
where the notation $(\,\cdot\,)^{\bullet}$ signifies quantities associated with a nominal AC power flow solution, and $b_{ij}$ on the right-hand side of (5a) is typically the cold-start version defined in (4). The bias $\bm{\gamma}^{hot}$ aims to capture the effects of branch losses on phase angles and $\bm{\rho}^{hot}$ incorporates the branch losses in the line flow expressions. A notable limitation of the hot-start approach is its reliance on a single, pre-determined nominal operating point, denoted by the $(\bullet)$ superscript in its defining equations. The accuracy of the hot-start parameters is therefore highest near that specific point and can degrade significantly as conditions deviate. This presents a challenge when a DC power flow approximation is needed for a wide range of operating conditions, as selecting hot-start parameters based on a single nominal operating point may yield insufficient accuracy.
II-B3 Optimized Parameters for DC Power Flow (DCPF) [taheri2023optimizing]
@@
Our recent work in [taheri2023optimizing] proposes an alternative for selecting the $\mathbf{b}$, $\bm{\gamma}$, and $\bm{\rho}$ parameters by solving a single-level optimization problem. Providing a conceptual foundation for the approach proposed in this paper, the algorithm in [taheri2023optimizing] also optimizes these parameters over a range of operation. However, there is a key distinction between the algorithm in [taheri2023optimizing] and the algorithm in this paper: the approach in [taheri2023optimizing] does not explicitly consider the impacts of the optimized parameters on the solution to a DC-OPF problem, but rather seeks to balance the DC power flow’s accuracy with respect to all operating points within a specified range. Conversely, our proposed algorithm in this paper leverages machine learning techniques for training models with differentiable optimization layers to tailor the DC power flow parameter values so that the DC optimal power flow solutions align with the AC-OPF solutions. As shown numerically in Section IV, this enables substantial accuracy improvements relative to [taheri2023optimizing] in settings where the optimized DC power flow parameters are used in optimization problems like DC-OPF. This represents a significant conceptual and technical advancement over the single-level problem in [taheri2023optimizing]. While the prior work focused on tuning parameters to match the line flow behaviors of the DC and AC power flow equations, the present work introduces a bilevel optimization framework. The objective is no longer to fit the intermediate equations, but rather to tune the parameters such that the final optimal solution of the DC-OPF problem aligns with that of the AC-OPF problem. This solution-focused methodology is a key contribution of this paper.
@@
We next present our algorithm for optimizing the coefficient ($\mathbf{b}$) and bias ($\bm{\gamma}$ and $\bm{\rho}$) parameters. Our algorithm seeks to match the generators’ active power outputs from the DC-OPF solution to corresponding values from the AC-OPF solution.
@@ III Parameter Optimization Algorithm
Our parameter optimization algorithm, depicted in Fig. 1, has offline and online phases. The offline phase computes DC-OPF parameters ($\mathbf{b}$, $\bm{\gamma}$, and $\bm{\rho}$) to achieve a high-fidelity approximation of the AC-OPF solution over a range of operational conditions. The online phase employs these parameters to efficiently solve DC-OPF problems in various applications. While our numerical results focus on DC-OPF problems, the optimized DC-OPF parameters are applicable to many computationally intensive optimization problems that incorporate DC power flow models, such as unit commitment [vanAckooij2018, knueven2020mixed], expansion planning [majidi2016stochastic], and grid reliability enhancement [Liu2010].
@@
Our algorithm minimizes a loss function based on the discrepancy between DC-OPF and AC-OPF generator active power setpoints over various load scenarios. To perform this minimization, we compute the loss function’s gradients with respect to the parameters $\mathbf{b}$, $\bm{\gamma}$, and $\bm{\rho}$ for use in a nonlinear optimization method. Based on our numerical tests comparing various optimization methods, TNC’s approach to approximating the Hessian matrix from gradient information best achieved scalable and effective parameter optimization.
@@ III-A Loss Function
We first introduce a loss function $\mathcal{L}$ that sums the squared two-norm discrepancies between the generators’ active power setpoints from the AC-OPF problems ($\mathbf{p}_{m}^{\text{g,AC}}$) and the DC-OPF problems ($\mathbf{p}_{m}^{\text{g}}$) across a specified set of load scenarios $\mathcal{M}$:
$\mathcal{L}(\mathbf{b},\bm{\gamma},\bm{\rho})$ $=\frac{1}{|\mathcal{G}|\times|\mathcal{M}|}\sum_{m\in\mathcal{M}}\|\mathbf{p}_{m}^{\text{g}}(\mathbf{b},\bm{\gamma},\bm{\rho})-\mathbf{p}_{m}^{\text{g,AC}}\|^{2}_{2},$ (6)
where the constant $\frac{1}{|\mathcal{G}|\times|\mathcal{M}|}$ normalizes this function based on the number of generators and load scenarios. As shown in (6), $\mathbf{p}_{m}^{\text{g}}$ (and thus $\mathcal{L}(\mathbf{b},\bm{\gamma},\bm{\rho})$) is a function of the coefficient parameters $\mathbf{b}$ and the bias parameters $\bm{\gamma}$ and $\bm{\rho}$. The selection of the training scenarios in the set $\mathcal{M}$ is an important application-dependent step. Scenarios should be chosen to represent the operating conditions under which the optimized parameters will be used. In our experiments, we generate scenarios by scaling nominal loads based on a normal distribution; however, in practice, utilities could use historical data or forecast scenarios to create a training set tailored to their system.
@@
$\operatorname*{minimize}_{\mathbf{b},\bm{\gamma},\bm{\rho}}\quad\mathcal{L}(\mathbf{b},\bm{\gamma},\bm{\rho}).$ (7)
Using the sensitivity analysis that we describe in the next subsection, we adjust the coefficient and bias parameters ($\mathbf{b}$, $\bm{\gamma}$, and $\bm{\rho}$) to optimize the DC-OPF problem’s parameters. As detailed in Model 4, this task takes the form of a bilevel optimization problem, with the upper-level fine-tuning parameters based on outcomes from lower-level DC-OPF problems under a variety of operational scenarios.
@@
$\operatorname*{minimize}_{\mathbf{b},\bm{\gamma},\bm{\rho}}\quad$ $\frac{1}{|\mathcal{G}|\times|\mathcal{M}|}\sum_{m\in\mathcal{M}}\|\mathbf{p}_{m}^{\text{g}}(\mathbf{b},\bm{\gamma},\bm{\rho})-\mathbf{p}^{\text{g},\text{AC}}_{m}\|^{2}_{2}$ (8a)
@@
$\operatorname*{minimize}_{\mathbf{p}^{\text{g}}_{m},\bm{\theta}_{m}}$ $\quad{\mathbf{p}^{\text{g}}_{m}}^{\top}\text{diag}(\mathbf{c}_{2})\mathbf{p}^{\text{g}}_{m}+\mathbf{c}^{\top}_{1}\mathbf{p}^{\text{g}}_{m}+\mathbf{1}^{\top}\mathbf{c}_{0}$ (8b)
@@
$\mathbf{p}^{\text{g}}_{m}-\textbf{p}^{\text{d}}_{m}-\bm{\gamma}=\mathbf{A}^{\top}\Big(\text{diag}(\mathbf{b})\mathbf{A}\bm{\theta}_{m}+\bm{\rho}\Big)$ (8c)
@@
    $\alpha_{2}$: Curvature condition constant, between $\alpha_{1}$ and 11
@@
6while $k\leq\textit{max\_iter}$ and $\|\nabla\mathcal{L}(\mathbf{x}_{k})\|>\epsilon$ do
@@ III-B Sensitivity Analysis of Coefficient and Bias Parameters
$\mathbf{g}^{b}=\frac{\partial\mathcal{L}}{\partial\mathbf{b}}=\frac{2}{|\mathcal{G}|\times|\mathcal{M}|}\sum_{m\in\mathcal{M}}\left.\Big(\frac{\partial\mathbf{p}^{\text{g}}}{\partial\mathbf{b}}\right|_{\mathbf{p}^{\text{g}}_{m}}\Big)^{\top}\Big(\mathbf{p}_{m}^{\text{g}}-\mathbf{p}_{m}^{\text{g},AC}\Big),$ (9a)
$\mathbf{g}^{\gamma}=\frac{\partial\mathcal{L}}{\partial\bm{\gamma}}=\frac{2}{|\mathcal{G}|\times|\mathcal{M}|}\sum_{m\in\mathcal{M}}\left.\Big(\frac{\partial\mathbf{p}^{\text{g}}}{\partial\bm{\gamma}}\right|_{\mathbf{p}^{\text{g}}_{m}}\Big)^{\top}\Big(\mathbf{p}_{m}^{\text{g}}-\mathbf{p}_{m}^{\text{g},AC}\Big),$ (9b)
$\mathbf{g}^{\rho}=\frac{\partial\mathcal{L}}{\partial\bm{\rho}}=\frac{2}{|\mathcal{G}|\times|\mathcal{M}|}\sum_{m\in\mathcal{M}}\left.\Big(\frac{\partial\mathbf{p}^{\text{g}}}{\partial\bm{\rho}}\right|_{\mathbf{p}^{\text{g}}_{m}}\Big)^{\top}\Big(\mathbf{p}_{m}^{\text{g}}-\mathbf{p}_{m}^{\text{g},AC}\Big).$ (9c)
The Jacobian matrices $\frac{\partial\mathbf{p}^{\text{g}}}{\partial\mathbf{b}}$, $\frac{\partial\mathbf{p}^{\text{g}}}{\partial\bm{\gamma}}$, and $\frac{\partial\mathbf{p}^{\text{g}}}{\partial\bm{\rho}}$ are central to the gradients in (9). Since $\mathbf{p}^{\text{g}}$ is the solution to the convex DC-OPF optimization problem, these derivatives are found by implicitly differentiating the Karush-Kuhn-Tucker (KKT) optimality conditions of the DC-OPF problem [fiacco1990nonlinear, amos2017optnet]. This is a key step that makes our bilevel approach tractable.
To perform these implicit derivative computations, we leverage techniques developed for embedding convex optimization layers within deep learning architectures; specifically, we utilize cvxpylayers[cvxpylayers2019], which allows for automatic differentiation through convex optimization problems specified in CVXPY. This integration facilitates efficient and scalable gradient computations necessary for our optimization framework. Our implementation defines the DC-OPF problem using CVXPY’s modeling syntax and uses cvxpylayers to create a differentiable layer. During the forward pass, this layer solves the DC-OPF problem to obtain $\mathbf{p}^{\text{g}}$. In the backward pass, it automatically computes the required Jacobians by differentiating the KKT system, leveraging the methods in [cvxpylayers2019, amos2017optnet].